All strictly monotonically decreasing sequences are convergent.

True or false
1
Expert's answer
2020-03-02T09:57:13-0500

False. A sequence "\{a_n\}_{n\in \mathbb{N}}" is called strictly monotonically decreasing if and only if "a_{n+1} < a_n" for all "n\in \mathbb{N}". Since "-(n+1) = -n-1 < -n", the sequence "\{-n\}_{n\in \mathbb{N}}" satisfies this definition and is strictly monotonically decreasing. Assume that it converges to some "a\in \mathbb{R}". Then there is "n_0\in \mathbb{N}" such that "a-1 < -m < a+1" for all "m\geq n_0". By the Archimedean property of the real numbers, there is "n_1\in \mathbb{N}" such that "-n_1 \leq a-1". Letting "n_2 = \operatorname{max}(n_0, n_1)", we have that "a-1 < -n_2" and "-n_2 \leq a-1", a contradiction. Therefore, "\{-n\}_{n\in \mathbb{N}}" is not convergent.